Let X denote the distance (m) that an animal moves from its birth site to the first territorial vacancy it encounters. Suppose that for banner-tailed kangaroo rats, X has an exponential distribution with parameter λ = 0.01457.

(a) What is the probability that the distance is at most 100 m? At most 200 m? Between 100 and 200 m? (Round your answers to four decimal places.)

at most 100 m     
at most 200 m     
between 100 and 200 m     


(b) What is the probability that distance exceeds the mean distance by more than 2 standard deviations? (Round your answer to four decimal places.)


(c) What is the value of the median distance? (Round your answer to two decimal places.)
Answer #1

The solution to this problem is given by

Solution Given Le rovi x has exponential aistribution parameta L. 1457 Xu Rap( h = 0.01457) -hu P(Xex) = i-e -0.01457x - l-e(6) Mear distance = M = I= 1 = 68.634 h 0.01457 Standard deviation == 1 = 1 0.01457 681634 P ( alistance encreeds mean distana) ku I-0.5 -0.5 e 11 -hau = ) la e hay ln (0.5) =) -han = -0.69314718) =124 + 0.69314718) 0.01457 21+4705735883 0,= + 47.57